How many liters of Cl2 gas are occupied by 710g of it at STP?
1
Expert's answer
2020-05-10T14:57:23-0400

T = 273.15 K

P = 101.325*103 Pa

m(Cl2) = 710g

R = 8.314 J/K*mol

PV = m*R*T/M

V = m*R*T/M*P = (710g*8.314 J/K*mol*273.15 K)/(71g/mol*101.325*103 Pa) = 0.224 m3 = 224 L